OpenStudy (anonymous):

Find the first six terms of the sequence. a1 = -8, an = 5 • an-1

OpenStudy (ybarrap):

Rewriting your equations: $$ a_1=-8\\ a_n=5a_{n-1} $$ Here, \(n\) represents the \(n^{th}\) term of the series and \(a_n\) represents the value of that term; so, \(a_1\) is the first term, \(a_2\) is the second term and \(a_6\) is the sixth term. So to get \(a_6\), the sixth term, all you need to do is set \(n=6\) in the second equation: $$ a_6=5a_{(6-1)}=5a_{5} $$ Which means you'll need \(a_5\) to solve: $$ a_5=5a_{(5-1)}=5a_{4} $$ Keep going like this until you get $$ a_2=5a_{(2-1)}=5a_{1} $$ Then you'll have everything you need to get \(a_6\). There is actually a simpler way, but for now you should do this to understand the series. Make sense?

OpenStudy (anonymous):

not really..